Solve and check. 1s+s-6/2/3=4/1/3. I get the answer 5.5, but how do you check it.
Answer by funmath(2933)

Solve and check. 1s+s-6 2/3=4 1/3.
s+s-6 2/3=4 1/3
2s-6 2/3+6 2/3=4 1/3+6 2/3
2s=10 3/3
2s=11
2s/2=11/2
s=11/2 or 5.5
:
Check by substituting your answer into the original equation where the variables are and see if both sides of the equation equal.
1(5.5)+(5.5)-6 2/3=4 1/3
5.5+5.5-6 2/3= 4 1/3
11-6 2/3=4 1/3
10 3/3-6 2/3=4 1/3
4 1/3= 4 1/3 You're right!!!
Happy Calculating!!!